OpenStudy (anonymous):

i need help and solving this problem (3-2i)-(x+yi)=(2i-3i)?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

sure it is (2i-3i)?

OpenStudy (anonymous):

because 2i - 3i = -i

OpenStudy (anonymous):

maybe 2 - 3i instead?

OpenStudy (anonymous):

is this the answer

OpenStudy (anonymous):

it says what are the values of x and y

OpenStudy (anonymous):

in this u need to compare the real and imaginary parts in both sides of the equation... so u hav 3-x=0=>x=3 and -i(2+y)=-i =>y=-2